### v1

@@ 4.5.1 Evaluation of Components
To thoroughly investigate the effectiveness of each proposed component and validate our design choices, we conduct comprehensive ablation studies on the Refer-KITTI. The detailed experimental results are presented in Table 5. We investigate the possibility of using a lightweight text encoder to reduce the computational complexity of our framework. Specifically, we replace the original RoBERTa encoder with MobileCLIP, a more parameter-efficient alternative. However, this modification leads to significant performance degradation across all metrics (HOTA decreases from 46.56 to 39.95, AssA drops from 57.33 to 52.32). These results indicate that while MobileCLIP offers advantages in terms of computational efficiency, the sophisticated language understanding capabilities provided by RoBERTa are essential for accurate referring expression comprehension in RMOT tasks. The substantial performance gap suggests that maintaining robust language representation power is crucial, even at the cost of higher computational requirements.
Further incorporating the Progressive Semantic-Decoupled Query Learning (PSDQL) module is analogous to the processing of the dorsal stream and ventral stream in human visual system. We observe consistent and substantial gains across all evaluation metrics (HOTA: 48.36, DetA: 39.97, AssA: 58.69). This comprehensive improvement suggests that explicit modeling of motion dynamics significantly enhances both detection and association capabilities.
Table 5: Ablation studies of different components in CDRMT. In the table, ‘w.’ indicates ‘with’ (component used) and ‘w/o.’ indicates ‘without’ (component not used).
[…]
Baseline RoBERTa 46.56 37.97 57.33 49.69 60.10 60.02 89.67 90.33 124.16 274.25
MobileCLIP 39.95 32.80 52.32 48.68 56.33 51.69 80.23 81.07 117.85 253.54
w.BIF RoBERTa 46.96 38.65 57.11 55.75 54.64 61.35 88.73 90.43 124.97 276.10
w.PSDQL RoBERTa 48.36 39.97 58.69 54.86 58.07 62.79 89.68 90.32 125.48 283.46
   w/o. ${Q}^{\prime}_{t}$ RoBERTa 48.02 39.45 58.62 55.46 56.41 63.09 88.77 90.52 125.02 280.91
   w/o. $\hat{Q}^{\prime}_{t}$ RoBERTa 47.89 39.63 58.04 54.66 57.64 61.94 89.77 90.60 124.97 280.36
w.SCC RoBERTa 47.82 39.05 58.75 56.98 54.08 63.55 88.14 90.26 124.83 275.63
CDRMT (ours) RoBERTa 49.35 40.34 60.56 54.54 59.30 64.70 89.80 90.61 126.00 287.03
To rigorously validate the necessity of our progressive design, we conduct controlled experiments by ablating each stream independently. When removing the static object query ${Q}^{\prime}_{t}$, which is responsible for encoding appearance and attribute information, the HOTA metric drops to 48.02. This degradation is particularly evident in detection accuracy (DetA decreases by 1.3%), indicating that the model’s ability to precisely locate objects based on their visual attributes is compromised. Similarly, ablating the spatial motion information query $\hat{Q}^{\prime}_{t}$ leads to a performance decline (HOTA: 47.89), with a more pronounced impact on association accuracy (AssA decreases by 1.1%). This suggests that without explicit motion state modeling, the model struggles to maintain consistent object identities across frames, especially in challenging scenarios with complex motion patterns or object interactions.
Figure 9: Qualitative comparison of tracking results across different model variants on challenging examples. From left to right: TransRMOT (baseline), CDRMT without spatial motion query processing ($\hat{Q}^{\prime}_{t}$), CDRMT without static object query processing (${Q}^{\prime}_{t}$), and our complete CDRMT model.
@@
Finally, we evaluate the contribution of the Structural Consensus Constraint (SCC) module, which serves as an additional supervisory signal to ensure comprehensive understanding of both static and dynamic features. The experimental results demonstrate that SCC not only helps maintain robust tracking performance but also enhances the model’s interpretability by enforcing semantic alignment between visual tracking results and language descriptions.
@@ 4.5.3 Weight of referring Threshold
4.5.4 Different Components of QSI Module
Table 8 presents a comprehensive ablation study of our Query Semantic Injection (QSI) module. Without QSI, the baseline model achieves performance (47.16 HOTA, 38.82 DetA, 57.47 AssA), indicating the limitations of conventional feature fusion approaches. When simultaneously injecting both static object and spatial motion information ($\hat{\mathbf{f}}_{so}$ $\hat{\mathbf{f}}_{sm}$), we observe initial improvements in all metrics (47.31 HOTA, 39.02 DetA, 57.82 AssA). However, this parallel injection strategy, while beneficial, fails to fully exploit the hierarchical nature of semantic information. In exploring semantic injection ordering, we find that injecting spatial motion information before static object attributes ($\hat{\mathbf{f}}_{sm}$rightarrow $\hat{\mathbf{f}}_{so}$) leads to suboptimal performance (47.63 HOTA), as the model struggles to establish stable object representations without first understanding basic object characteristics. Conversely, following the cognitive process of human vision by first injecting static object features followed by motion information ($\hat{\mathbf{f}}_{so}$rightarrow $\hat{\mathbf{f}}_{sm}$) yields superior results (48.06 HOTA), demonstrating the importance of progressive semantic understanding from basic to complex features. The impact of query-specific injection reveals interesting patterns. Detection query injection ($Q^{Det}_{inject}$) significantly enhances localization accuracy (40.22 DetA) by providing explicit semantic guidance for object detection. Meanwhile, track query injection ($Q^{Tra}_{inject}$) exhibits stronger temporal modeling capabilities, as evidenced by improved association scores (60.17 AssA). The synergistic combination of both injection strategies ($Q^{Det}_{inject}$ $Q^{Tra}_{inject}$) achieves optimal performance across all metrics (49.35 HOTA, 40.34 DetA, 60.56 AssA), validating our hypothesis that comprehensive semantic understanding requires fine-grained semantic guidance.
@@ 4.5.4 Different Components of QSI Module
4.5.5 Different Components of SCC Module
Table 9: Ablation study on Structural Consensus Constraint (SCC) module.
@@ 4.5.5 Different Components of SCC Module
Table 9 presents the impact of different Structural Consensus Constraint (SCC) configurations on model performance. The baseline model (“-”) without any consistency constraint performs worst across all metrics. The traditional Point-wise Consistency (PwC), which only operates at the single sample level by directly comparing the original text embedding $\mathbf{e}$ with its corresponding reconstructed text embedding $\mathbf{e}^{\prime}$, shows limited improvement (HOTA increases by only 0.16%). In contrast, our proposed structural consistency approach includes two key components. The Distance Consistency (DC) constraint leverages relationships between three different text embeddings (static object embedding $\mathbf{e}_{so}$, spatial motion embedding $\mathbf{e}_{sm}$, and global sentence representation $\mathbf{e}_{g}$) and their reconstructed counterparts. By maintaining the relative distance structure between these embedding pairs rather than direct one-to-one mapping, it achieves significant improvement (HOTA reaches 48.41). Similarly, the Angle Consistency (AC) constraint considers angular relationships formed by triplets of embeddings, capturing higher-order semantic structures and achieving a HOTA of 48.19. The most significant performance gains come from combining both constraints into a complete Structural Consistency Constraint (SCC). When $\lambda_{angle}=0.4$ , the model achieves optimal performance across all metrics. This demonstrates the strong complementarity between the two constraints: distance consistency preserves semantic similarities, while angle consistency captures more complex structural relationships between embeddings.
@@ 5 Discussion
Overall, our Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework demonstrates significant advantages in handling complex semantic information in RMOT tasks. By mimicking the dual-stream processing mechanism of human visual system, our approach effectively disentangles static object attributes from spatial motion information, enabling more precise object localization and tracking. However, we acknowledge certain limitations of our current approach, particularly in challenging environmental conditions. As shown in Figure 6 #273, our model occasionally struggles with appearance-based discrimination under extreme lighting conditions, mistakenly identifying a white car as silver due to illumination variations. Additionally, we observe that the association metrics still fall short compared to some state-of-the-art methods, indicating room for improvement in maintaining consistent object identities across frames.
Looking forward, these observations motivate several important directions for future research. We plan to construct a dedicated RMOT dataset focusing on challenging scenarios to better address the limitations in extreme conditions. Furthermore, improving the association capability while maintaining strong detection performance remains a crucial direction for future work. We believe these developments will contribute to more robust and reliable RMOT systems capable of handling a broader range of real-world scenarios.
@@ 6 Conclusion
In this paper, we present a novel Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework that effectively addresses the challenges of processing complex semantic information in multi-source information fusion. Drawing inspiration from the dual-stream processing mechanism in human visual system, our approach explicitly disentangles static object attributes and spatial motion information in referring expressions, enabling more precise and robust tracking results. The proposed Bidirectional Interactive Fusion module enables structured cross-modal interactions while preserving modality-specific characteristics. Additionally, the Progressive Semantic-Decoupled Query Learning mechanism hierarchically injects these two types of information into different decoder layers, while the Structural Consensus Constraint ensures comprehensive semantic alignment between visual and linguistic modalities. Extensive experiments on multiple benchmark datasets demonstrate that CDRMT achieves SOTA performance, which validates the effectiveness of our method. Future research could explore the extension of our cognitive disentanglement approach to other information fusion tasks and investigate more sophisticated mechanisms for modeling the dynamic interplay between different semantic components in natural language descriptions.

### v3

@@ 4.5.1 Evaluation of Components
To thoroughly investigate the effectiveness of each proposed component and validate our design choices, we conduct comprehensive ablation studies on the Refer-KITTI. The detailed experimental results are presented in Table 5. The baseline architecture (row ①), as our comparison, we first examine the effect of integrating the Bidirectional Interactive Fusion (BIF) module (row ②). This integration yields quantifiable improvements in holistic tracking accuracy (HOTA increases by 0.9%) and detection performance (DetA improves by 1.8%). These enhancements empirically validate our hypothesis that structured bidirectional cross-modal interactions prior to the encoder stage facilitate more effective modality-specific feature refinement while preserving representational integrity. Further incorporating the Progressive Semantic-Decoupled Query Learning (PSDQL) module (row ③) is analogous to the processing of the dorsal stream and ventral stream in human visual system. We observe consistent and substantial gains across all evaluation metrics (HOTA: 48.36, DetA: 39.97, AssA: 58.69). To rigorously validate the necessity of our progressive design, we conduct controlled experiments by ablating each stream independently. When removing the static object query ${Q}^{\prime}_{t}$ (row ④), which is responsible for encoding appearance and attribute information, the HOTA metric drops to 48.02. This degradation is particularly evident in detection accuracy (DetA decreases by 1.3%), indicating that the model’s ability to precisely locate objects based on their visual attributes is compromised. Similarly, ablating the spatial motion information query $\hat{Q}^{\prime}_{t}$ (row ⑤) leads to a performance decline (HOTA: 47.89), with a more pronounced impact on association accuracy (AssA decreases by 1.1%). This suggests that without explicit motion state modeling, the model struggles to maintain consistent object identities across frames, especially in challenging scenarios with complex motion patterns or object interactions.
Figure 10: Qualitative comparison of tracking results across different model variants on challenging examples. From left to right: TransRMOT (baseline), CDRMT without spatial motion query processing ($\hat{Q}^{\prime}_{t}$), CDRMT without static object query processing (${Q}^{\prime}_{t}$), and our CDRMT model.
@@
We further investigate the interaction effects between multiple components to evaluate their complementarity. The joint implementation of BIF and PSDQL (row ⑥) yields performance superior to either component in isolation (HOTA: 48.67, DetA: 40.05, AssA: 59.31). The combination of BIF and Structural Consensus Constraint (SCC) (row ⑦) demonstrates moderate improvement over the BIF-only configuration, while the integration of PSDQL with SCC (row ⑧) produces substantial performance gains (HOTA: 49.06). The experimental results demonstrate that SCC not only helps maintain robust tracking performance but also enhances the model’s interpretability by enforcing semantic alignment between visual tracking results and language descriptions.
@@ 4.5.3 Weight of Referring Threshold
Table 9: Ablation study on Structural Consistency Constraint (SCC) module.
@@
4.5.4 Different Components of QSI Module
Table 8 presents a comprehensive ablation study of our Query Semantic Injection (QSI) module. Without QSI, the model achieves performance (47.16 HOTA, 38.82 DetA, 57.47 AssA), indicating the limitations of conventional feature fusion approaches. When simultaneously injecting both static object and spatial motion information ($\hat{\mathbf{f}}_{so}$ $\hat{\mathbf{f}}_{sm}$), we observe initial improvements in all metrics (47.31 HOTA, 39.02 DetA, 57.82 AssA). However, this parallel injection strategy, while beneficial, fails to fully exploit the hierarchical nature of semantic information. In exploring semantic injection ordering, we find that injecting spatial motion information before static object attributes ($\hat{\mathbf{f}}_{sm}$rightarrow $\hat{\mathbf{f}}_{so}$) leads to suboptimal performance (47.63 HOTA), as the model struggles to establish stable object representations without first understanding basic object characteristics. Conversely, following the cognitive process of human vision by first injecting static object features followed by motion information ($\hat{\mathbf{f}}_{so}$rightarrow $\hat{\mathbf{f}}_{sm}$) yields superior results (48.06 HOTA), demonstrating the importance of progressive semantic understanding from basic to complex features. The impact of query-specific injection reveals interesting patterns. Detection query injection ($Q^{Det}_{inject}$) significantly enhances localization accuracy (40.22 DetA) by providing explicit semantic guidance for object detection. Meanwhile, track query injection ($Q^{Tra}_{inject}$) exhibits stronger temporal modeling capabilities, as evidenced by improved association scores (60.17 AssA). The synergistic combination of both injection strategies ($Q^{Det}_{inject}$ $Q^{Tra}_{inject}$) achieves optimal performance across all metrics (49.35 HOTA, 40.34 DetA, 60.56 AssA), validating our hypothesis that comprehensive semantic understanding requires fine-grained semantic guidance.
4.5.5 Different Components of SCC Module
@@
Table 9 presents the impact of different Structural Consistency Constraint (SCC) configurations on model performance. The baseline model (“-”) without any consistency constraint performs worst across all metrics. The traditional Point-wise Consistency (PwC), which only operates at the single sample level by directly comparing the original text embedding $\mathbf{e}$ with its corresponding reconstructed text embedding $\mathbf{e}^{\prime}$, shows limited improvement (HOTA increases by only 0.16%). In contrast, our proposed structural consistency approach includes two key components. The Distance Consistency (DC) constraint leverages relationships between three different text embeddings (static object embedding $\mathbf{e}_{so}$, spatial motion embedding $\mathbf{e}_{sm}$, and global sentence representation $\mathbf{e}_{g}$) and their reconstructed counterparts. By maintaining the relative distance structure between these embedding pairs rather than direct one-to-one mapping, it achieves significant improvement (HOTA reaches 48.41). Similarly, the Angle Consistency (AC) constraint considers angular relationships formed by triplets of embeddings, capturing higher-order semantic structures and achieving a HOTA of 48.19. The most significant performance gains come from combining both constraints into a complete Structural Consistency Constraint (SCC). When $\lambda_{angle}=0.4$ , the model achieves optimal performance across all metrics. This demonstrates the strong complementarity between the two constraints: distance consistency preserves semantic similarities, while angle consistency captures more complex structural relationships between embeddings.
Figure 11: Visualization of heatmap from our CDRMT framework compared with baseline TransRMOT. The color intensity represents attention activation strength, with warmer colors (red and yellow) indicating regions of higher feature activation and cooler colors (blue) indicating regions of lower activation.
5 Visualization
To illustrate the effectiveness of our proposed CDRMT framework, we visualize the attention heatmaps from the final encoder layer in Figure 11. The integration of our method results in significantly more focused attention on regions that correspond to the referring text. As shown in Figure 11(a) with the expression "black cars in the left," while both models show attention on a white car positioned on the left, our model demonstrates heightened attention specifically towards the upper part of the vehicle where the roof is black, indicating a more nuanced understanding of the object’s attributes. This more targeted attention distribution is semantically aligned with the "black cars" description, even when the car body is predominantly white. Similarly, in Figure 11(b) with "cars in light color," the baseline model exhibits a more uniform attention distribution across the scene, whereas our approach shows concentrated attention specifically on vehicles with lighter color tones, demonstrating improved semantic discrimination. In Figure 11(c), our model exhibits highly discriminative attention patterns that concentrate exclusively on walking pedestrians on the left side of the scene, while successfully suppressing attention on both stationary pedestrians.
@@ 6 Discussion
Overall, our Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework demonstrates significant advantages in handling complex semantic information in RMOT tasks. By inspired the dual-stream processing mechanism of the human visual system, our approach effectively disentangles static object attributes from spatial motion information, enabling more precise object localization and tracking.
6.1 Limitations
However, we acknowledge certain limitations of our current approach, particularly in challenging environmental conditions. We have conducted further analysis of failure cases to better understand the limitations of our approach. Figure 12 illustrates several representative failure scenarios where CDRMT encounters difficulties. These include cases with extreme lighting conditions, severe occlusions, and low-resolution scenarios that challenge our cognitive disentanglement approach. Environmental variations impact performance, particularly with color-based attributes. For instance, our model misidentifies a white car as "silver" due to strong sunlight reflections (Figure 12 #178, #273). Crowded scenarios with occlusions present challenges, as demonstrated by our model’s failure to detect certain instances of "woman" in an outdoor cafe (Figure 12 #201, #364) and in a busy street scene (#364). Low-resolution scenarios with similar objects lead to misidentifications(Figure 12 #108, #295), where women with shorter hairstyles appear visually similar to men from a distance.
Our cognitive disentanglement approach effectively handles complex expressions by separating "what" and "where" pathways, but it also depends on visual feature quality. When these features become ambiguous due to resolution limitations or inherent visual similarities, even our enhanced semantic processing may fail to correctly associate with linguistic elements.
Figure 12: The failure cases of CDRMT.
6.2 Future Work
Looking forward, the limitations identified in our study motivate several important directions for future research. We plan to construct a dedicated RMOT dataset focusing on challenging scenarios to better address the limitations in extreme conditions. This dataset would specifically include scenes with varying lighting conditions, partial occlusions, and complex background environments to push the boundaries of RMOT research. In addition, the current dataset expression is relatively fixed and cannot achieve dynamic and fine-grained descriptions that change over time. We hope to solve this problem in the future. Exploring multi-granularity feature integration techniques could address the challenge of similar objects in low-resolution scenarios. The poor performance on rare motion patterns like "turning" highlights the need for specialized techniques to handle long-tailed distributions in motion expressions. It is worth mentioning that future work could explore integrating more advanced language models with stronger contextual reasoning capabilities. This would help the system better understand implicit spatial relationships and make more human-like inferences about object relationships and behaviors.
We believe these developments will contribute to more robust and reliable RMOT systems capable of handling a broader range of real-world scenarios, ultimately bringing us closer to human-level performance in language-guided multi-object tracking tasks.
@@ 7 Conclusion
In this paper, we present a novel Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework that effectively addresses the challenges of processing complex semantic information in multi-source information fusion. Drawing inspiration from the dual-stream processing mechanism in human visual system, our approach explicitly disentangles static object attributes and spatial motion information in referring expressions, enabling more precise and robust tracking results. The proposed Bidirectional Interactive Fusion module enables structured cross-modal interactions while preserving modality-specific characteristics. Additionally, the Progressive Semantic-Decoupled Query Learning mechanism hierarchically injects these two types of information into different decoder layers, while the Structural Consistency Constraint ensures comprehensive semantic alignment between visual and linguistic modalities. Extensive experiments on multiple benchmark datasets demonstrate that CDRMT achieves SOTA performance, which validates the effectiveness of our method. Future research could explore the extension of our cognitive disentanglement approach to other information fusion tasks and investigate more sophisticated mechanisms for modeling the dynamic interplay between different semantic components in natural language descriptions.